### v2

# LlamaFactory: Unified Efficient Fine-Tuning of 100+ Language Models

Abstract
Efficient fine-tuning is vital for adapting large language models (LLMs) to downstream tasks. However, it requires non-trivial efforts to implement these methods on different models. We present LlamaFactory, a unified framework that integrates a suite of cutting-edge efficient training methods. It allows users to flexibly customize the fine-tuning of 100+ LLMs without the need for coding through the built-in web UI LlamaBoard. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks. It has been released at https://github.com/hiyouga/LLaMA-Factory and already received over 13,000 stars and 1,600 forks.
LlamaFactory: Unified Efficient Fine-Tuning of 100+ Language Models
Yaowei Zheng1, Richong Zhang1, Junhao Zhang1, Yanhan Ye1, Zheyan Luo1, Yongqiang Ma2 1School of Computer Science and Engineering, Beihang University, China 2School of Software and Microelectronics, Peking University, China hiyouga@buaa.edu.cn, zhangrc@act.buaa.edu.cn, {zhang.jh,yeyanhan,akamya}@buaa.edu.cn, codingma@pku.edu.cn Demonstration video: https://youtu.be/W29FgeZEpus
1 Introduction
Large language models (LLMs) (Zhao et al., 2023) present remarkable reasoning capabilities and empower a wide range of applications, such as question answering (Jiang et al., 2023b), machine translation (Wang et al., 2023c; Jiao et al., 2023a), and information extraction (Jiao et al., 2023b). Subsequently, a substantial number of LLMs are developed and accessible through open-source communities. For example, Hugging Face’s open LLM leaderboard (Beeching et al., 2023) boasts over 5,000 models, offering convenience for individuals seeking to leverage the power of LLMs.
Fine-tuning extremely large number of parameters with limited resources becomes the main challenge of adapting LLM to downstream tasks. A popular solution is efficient fine-tuning (Houlsby et al., 2019; Hu et al., 2022; Ben Zaken et al., 2022; Dettmers et al., 2023; Zhao et al., 2024), which reduces the training cost of LLMs when adapting to various tasks. However, the community contributes various methods for efficient fine-tuning LLMs, lacking a systematic framework that adapts and unifies these methods to different LLMs and provides a friendly interface for user customization.
To address the above problems, we develop LlamaFactory, a framework that democratizes the fine-tuning of LLMs. It unifies a variety of efficient fine-tuning methods through scalable modules, enabling the fine-tuning of hundreds of LLMs with minimal resources and high throughput. In addition, it streamlines commonly used training approaches, including generative pre-training (Radford et al., 2018), supervised fine-tuning (SFT) (Wei et al., 2022), reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), and direct preference optimization (DPO) Rafailov et al. (2023). Users can leverage command-line or web interfaces to customize and fine-tune their LLMs with minimal or no coding effort.
LlamaFactory consists of three main modules: Model Loader, Data Worker and Trainer. We minimize the dependencies of these modules on specific models and datasets, allowing the framework to flexibly scale to hundreds of models and datasets. Concretely, we first establish a model registry where the Model Loader can precisely attach adapters to the pre-trained models by identifying exact layers. Then we develop a data description specification that allows the Data Worker to gather datasets by aligning corresponding columns. Furthermore, we provide plug-and-play implementations of efficient fine-tuning methods that enable the Trainer to active by replacing default ones. Our design allows these modules to be reused across different training approaches, significantly reducing integration cost of new methods.
LlamaFactory is implemented with PyTorch (Paszke et al., 2019) and significantly benefits from open-source libraries, such as Transformers (Wolf et al., 2020), PEFT (Mangrulkar et al., 2022), and TRL (von Werra et al., 2020). On the basis, we provide an out-of-the-box framework with a higher level of abstraction. Additionally, we build LlamaBoard with Gradio (Abid et al., 2019), enabling fine-tuning LLMs with no coding efforts required.
LlamaFactory is open-sourced under the Apache-2.0 license. It has already garnered over 13,000 stars and 1,600 forks on the GitHub [^1] https://github.com/hiyouga/LLaMA-Factory, and hundreds of open-source models have been built upon LlamaFactory on the Hugging Face Hub [^2] https://huggingface.co/models?other=llama-factory. For example, the well-known GemSUra-7B (Nguyen et al., 2024) is built based on LlamaFactory, which first reveals the cross-lingual abilities of Gemma (Mesnard et al., 2024). Furthermore, dozens of research studies utilize our framework to explore new methods in LLMs, such as Wang et al. (2023a); Yu et al. (2023); Bhardwaj et al. (2024).
2 Efficient Fine-Tuning Techniques
Efficient LLM fine-tuning techniques can be divided into two main categories: those focused on optimization and those aimed at computation. The primary objective of efficient optimization techniques is to adjust the parameters of LLMs while keeping costs to a minimum. On the other hand, efficient computation methods seek to decrease the time or space for the required computation in LLMs. The methods included in LlamaFactory are listed in Table 1. We will present these efficient fine-tuning techniques and show the substantial efficiency improvement achieved by incorporating them into our framework in the following sections.
Table 1: Efficient fine-tuning techniques featured in LlamaFactory. Techniques that are compatible with each other are marked with ✓, while those that are not compatible are marked with ✗.
Freeze-tuning GaLore LoRA DoRA
Mixed precision
Checkpointing
Flash attention
S2 attention
Quantization
Unsloth
2.1 Efficient Optimization
Firstly, we provide an overview of the efficient optimization techniques utilized in LlamaFactory. The freeze-tuning method (Houlsby et al., 2019) involves freezing a majority of parameters while fine-tuning the remaining parameters in a small subset of decoder layers. Another method called gradient low-rank projection (GaLore) (Zhao et al., 2024) projects gradients into a lower-dimensional space, facilitating full-parameter learning in a memory-efficient manner. On the contrary, the low-rank adaptation (LoRA) (Hu et al., 2022) method freezes all pre-trained weights and introduces a pair of trainable low-rank matrices to the designated layer. When combined with quantization, this approach is referred to as QLoRA (Dettmers et al., 2023), which additionally reduces the memory usage. The weight-decomposed low-rank adaptation (DoRA) (Liu et al., 2024) method breaks down pre-trained weights into magnitude and direction components, applying LoRA solely on the directional components to enhance the fine-tuning of LLMs. LoRA+ (Hayou et al., 2024) is proposed to overcome the sub-optimality of LoRA.
2.2 Efficient Computation
In LlamaFactory, we integrate a range of efficient computational techniques. Commonly utilized techniques encompass mixed precision training (Micikevicius et al., 2018) and activation checkpointing (Chen et al., 2016). Drawing insights from the examination of the input-output (IO) expenses of the attention layer, flash attention (Dao et al., 2022) introduces a hardware-friendly approach to enhance attention computation. S2 attention (Chen et al., 2024b) tackles the challenge of extending context in block-sparse attention, thereby diminishing memory usage in fine-tuning long-context LLMs. Various quantization strategies (Dettmers et al., 2022a; Frantar et al., 2023; Lin et al., 2023; Egiazarian et al., 2024) decrease memory requirements in large language models (LLMs) by utilizing lower-precision representations for weights. Nevertheless, the fine-tuning of quantized models is restricted to the adapter-based techniques like LoRA (Hu et al., 2022). Unsloth (Han and Han, 2023) incorporates Triton (Tillet et al., 2019) for implementing the backward propagation of LoRA, which reduces floating-point operations (FLOPs) during gradient descent and leads to expedited LoRA training.
LlamaFactory effectively combines these techniques into a cohesive structure that greatly enhances the efficiency of LLM fine-tuning. This results in a reduction of the memory footprint from 18 bytes per parameter during mixed precision training (Micikevicius et al., 2018) or 8 bytes per parameter in bfloat16 training (Le Scao et al., 2022) to only 0.6 bytes per parameter. Further elaboration on our framework will be provided in the subsequent section.
3 LlamaFactory Framework
LlamaFactory consists of three main modules: Model Loader, Data Worker, and Trainer. The Model Loader prepares various architectures for fine-tuning, which supports over 100 LLMs. The Data Worker processes data from different tasks through a well-designed pipeline, which support more than 50 datasets. The Trainer unifies efficient fine-tuning methods to adapt these models on different tasks and datasets, which offers four training approaches. LlamaBoard provides a friendly visual interface for the above modules, enabling users to configure and launch individual LLM fine-tuning process in a codeless way and monitor the training status on the fly. We illustrate the overall architecture of LlamaFactory in Figure 1.
LlamaBoardArgument ConfiguratorTraining Status MonitorTrainerOptimizationApproachesLoRA+GaLoreOffloadPartitionPre-trainSFTRLHFDPOModel LoaderData WorkerInitializationPatchesQuantizationAdaptersAligningLoadingPreprocessMerging100+ Pre-Trained Models50+ NLP Datasets
Figure 1: The architecture of LlamaFactory.
3.1 Model Loader
This section initially presents the four components in Model Loader: Model Initialization, Model Patching, Model Quantization, and Adapter Attaching, followed by a description of our approach of adapting to a wide range of devices by handling the parameter float-point precision when fine-tuning.
Model Initialization
We employ the AutoModel API of Transformers (Wolf et al., 2020) for loading models and initializing parameters. In order to make our framework compatible with different model architectures, we establish a model registry to store the type of each layer, thereby facilitating the utilization of efficient fine-tuning techniques more straightforward. In cases where the vocabulary size of the tokenizer exceeds the capacity of the embedding layer, we resize the layer and initialize new parameters with noisy mean initialization. To determine the scaling factor for RoPE scaling (Chen et al., 2023), we compute it as the ratio of the maximum input sequence length to the context length of the model.
Model Patching
To enable flash attention and S2 attention, we employ a monkey patch to replace the forward computation of models. Nevertheless, since flash attention has been supported since Transformers 4.34.0, we use the API to enable flash attention. To prevent excessive partitioning of the dynamic modules, we set the mixture-of-experts (MoE) blocks as leaf modules when we optimize it under DeepSpeed ZeRO-3 (Rasley et al., 2020).
Model Quantization
Dynamically quantizing models to 8 bits or 4 bits with LLM.int8 (Dettmers et al., 2022a) can be performed through the bitsandbytes library (Dettmers, 2021). For 4-bit quantization, we utilize the double quantization and 4-bit normal float as QLoRA (Dettmers et al., 2023). We also support fine-tuning the models quantized by the post-training quantization (PTQ) methods, including GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2023), and AQLM (Egiazarian et al., 2024). Note that we cannot directly fine-tune the quantized weights; thus, the quantized models are only compatible with adapter-based methods.
Adapter Attaching
We automatically identify the appropriate layers using the model registry to attach adapters. Adapters are attached to a subset of layers by default to save memory, but attaching them to all linear layers may yield better performance (Dettmers et al., 2023). The PEFT (Mangrulkar et al., 2022) library offers an extremely convenient way to attach adapters such as LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), and DoRA (Liu et al., 2024). We replace the backward computation with the one of Unsloth (Han and Han, 2023) to accelerate LoRA. To perform reinforcement learning from human feedback (RLHF), we add a value head upon the model, a linear layer that maps the representation of each token to a scalar.
Precision Adaptation
We handle the floating-point precision of pre-trained models based on the capabilities of devices. For NVIDIA GPUs, we use bfloat16 precision if the computation capability is 8.0 or higher. Otherwise, float16 is adopted. We use float16 for Ascend NPUs and AMD GPUs and float32 for non-CUDA devices. Note that loading bfloat16 models with float16 precision may lead to overflow issues. In mixed precision training, we set all trainable parameters to float32. Nevertheless, we retain the trainable parameters as bfloat16 in bfloat16 training.
3.2 Data Worker
We develop a data processing pipeline, including dataset loading, dataset aligning, dataset merging and dataset pre-processing. It standardizes datasets of different tasks into a unified format, enabling us to fine-tune models on datasets in various formats.
Dataset Loading
We utilize the datasets (Lhoest et al., 2021) library to load the data, which allows the users to load remote datasets from the Hugging Face Hub or read local datasets via scripts or through files. The datasets library significantly reduces memory overhead during data processing and accelerates sample querying using Arrow (Apache, 2016). By default, the whole dataset is downloaded to local disk. However, if a dataset is too large to be stored, our framework provides dataset streaming to iterate over it without downloading.
Table 2: Dataset structures in LlamaFactory.
Plain text [{"text": "…"}, {"text": "…"}]
Alpaca-like data [{"instruction": "…", "input": "…", "output": "…"}]
ShareGPT-like data [{"conversations": [{"from": "human", "value": "…"}, {"from": "gpt", "value": "…"}]}]
Preference data [{"instruction": "…", "input": "…", "output": ["…", "…"]}]
Standardized data {"prompt": [{"role": "…", "content": "…"}],
"response": [{"role": "…", "content": "…"}],
"system": "…", "tools": "…"}
Dataset Aligning
To unify the dataset format, we design a data description specification to characterize the structure of datasets. For example, the alpaca dataset has three columns: instruction, input and output (Taori et al., 2023). We convert the dataset into a standard structure that is compatible with various tasks according to the data description specification. Some examples of dataset structures are shown in Table 2.
Dataset Merging
The unified dataset structure provides an efficient approach for merging multiple datasets. For the datasets in non-streaming mode, we simply concatenate them before the datasets are shuffled during training. However, in streaming mode, simply concatenating the datasets impedes data shuffling. Therefore, we offer methods to alternately read the data from different datasets.
Dataset Pre-processing
LlamaFactory is designed for fine-tuning the text generative models, which is primarily used in chat completion. Chat template is a crucial component in these models, because it is highly related to the instruction-following abilities of these models. Therefore, we provide dozens of chat templates that can be automatically chosen according to the model type. We encode the sentence after applying the chat template using the tokenizer. By default, we only compute loss on the completions, while the prompts are disregarded (Taori et al., 2023). Optionally, we can utilize sequence packing (Krell et al., 2021) to reduce the training time, which is automatically enabled when performing generative pre-training. Appendix C shows details of our template design.
3.3 Trainer
Efficient Training
We integrate state-of-the-art efficient fine-tuning methods, including LoRA+ (Hayou et al., 2024) and GaLore (Zhao et al., 2024) to the trainer by replacing the default component. These training approaches are independent of the trainer, making them easily applicable to various tasks. We utilize the trainer of Transformers (Wolf et al., 2020) for pre-training and SFT, while adopting the trainers of TRL (von Werra et al., 2020) for RLHF and DPO. The tailored data collators are leveraged to differentiate trainers of various training approaches. To match the input format of the trainer for preference data, we build $2n$ samples in a batch where the first $n$ samples are chosen examples and the last $n$ samples are rejected examples.
Model-Sharing RLHF
Allowing RLHF training on consumer devices is a useful property for LLM fine-tuning. However, it is difficult because RLHF training requires four different models. To address this problem, we propose model-sharing RLHF, enabling entire RLHF training with no more than one pre-trained model. Concretely, we first train an adapter and a value head with the objective function for reward modeling, allowing the model to compute reward scores. Then we initialize another adapter and value head and train them with the PPO algorithm (Ouyang et al., 2022). The adapters and value heads are dynamically switched through the set_adapter and disable_adapter APIs of PEFT (Mangrulkar et al., 2022) during training, allowing the pre-trained model to serve as policy, value, reference, and reward models simultaneously. To the best of our knowledge, this is the first method that supports RLHF training on consumer devices.
Distributed Training
We can combine the above trainers with DeepSpeed (Rasley et al., 2020) for distributed training. Leveraging the DeepSpeed ZeRO optimizer, the memory consumption can be further reduced via partitioning or offloading.
3.4 Utilities
Accelerated Inference
During inference time, we reuses the chat template from the data worker to build the model inputs. We offer support for sampling the model outputs using Transformers (Wolf et al., 2020) and vLLM (Kwon et al., 2023), both of which support stream decoding. Additionally, we implement an OpenAI-style API that utilizes the asynchronous LLM engine and paged attention of vLLM to provide high-throughput concurrent inference services, facilitating the deployment of fine-tuned LLMs into various applications.
Comprehensive Evaluation
We include several metrics for evaluating LLMs, including multiple-choice tasks such as MMLU (Hendrycks et al., 2021), CMMLU (Li et al., 2023a), and C-Eval (Huang et al., 2023), as well as calculating text similarity scores like BLEU-4 (Papineni et al., 2002) and ROUGE (Lin, 2004).
3.5 LlamaBoard: A Unified Interface for LlamaFactory
LlamaBoard is a unified user interface based on Gradio (Abid et al., 2019) that allows users to customize the fine-tuning of LLMs without writing any code. It offers a streamlined model fine-tuning and inference service, enabling users to easily leveraging 100+ LLMs and 50+ datasets in their practice. LlamaBoard has the following notable features:
Easy Configuration LlamaBoard allows users to customize fine-tuning arguments through interaction with the web interface. We provide default values for many arguments that are recommended for most users, simplifying the configuration process. Moreover, users can preview datasets on the web UI to check their customized format.
Monitorable Training During the training process, the training logs and loss curves are visualized and updated in real time, allowing users to monitor the training progress. This feature provides valuable insights to analyze the fine-tuning process.
Flexible Evaluation LlamaBoard supports calculating the text similarity scores on the datasets to automatically evaluate the model or performing human evaluation by chatting with the model.
Multilingual Support LlamaBoard provides localization files, facilitating the integration of new languages for rendering the interface. Currently we support three languages: English, Russian and Chinese, which allows a broader range of users to utilize LlamaBoard for fine-tuning LLMs.
4 Empirical Study
Table 3: Comparison of the training efficiency using different fine-tuning methods in LlamaFactory. The best result among GaLore, LoRA and QLoRA of each model is in bold.
Gemma-2B Llama2-7B Llama2-13B
Method Memory Throughput PPL Memory Throughput PPL Memory Throughput PPL
(GB) (Tokens/s) (GB) (Tokens/s) (GB) (Tokens/s)
Baseline / / 11.83 / / 7.53 / / 6.66
Full-tuning 17.06 3090.42 10.34 38.72 1334.72 5.56 / / /
Freeze-tuning 8.10 5608.49 11.33 15.69 2904.98 6.59 29.02 1841.46 6.56
GaLore 10.16 2483.05 10.38 15.43 1583.77 5.88 28.91 956.39 5.72
LoRA 7.91 3521.05 10.19 16.32 1954.07 5.81 30.09 1468.19 5.75
QLoRA 5.21 3158.59 10.46 7.52 1579.16 5.91 12.61 973.53 5.81
Table 4: Comparison of the performance (in terms of ROUGE) on specific tasks using different fine-tuning methods in LlamaFactory. The best result of each model is underlined, and the best result of each task is in bold.
CNN / DM XSum AdGen
Model Baseline FT GaLore LoRA QLoRA Baseline FT GaLore LoRA QLoRA Baseline FT GaLore LoRA QLoRA
LLaMA2-7B 12.94 22.87 22.40 22.70 22.61 13.89 27.69 27.64 28.80 28.05 0.61 20.51 19.61 20.29 20.45
Mistral-7B 14.39 22.03 22.99 23.47 23.28 15.87 23.57 28.00 30.41 30.44 7.82 20.14 20.90 20.99 20.56
Gemma-7B 15.97 22.07 / 22.41 22.44 15.31 25.13 / 28.67 29.02 11.57 19.99 / 20.62 19.81
Qwen1.5-7B 15.40 22.46 21.76 22.71 22.52 19.27 26.68 26.64 27.77 27.60 14.49 20.42 21.08 21.31 21.34
Yi-6B 16.85 22.40 22.68 22.98 22.97 18.24 27.09 28.25 28.71 29.21 13.34 19.68 20.06 20.97 20.31
ChatGLM3-6B 18.51 22.00 22.16 21.68 21.70 16.14 26.25 26.34 26.50 26.78 14.53 19.91 20.57 20.47 20.49
We systematically evaluate LlamaFactory from two perspectives: 1) the training efficiency in terms of memory usage, throughput and perplexity. 2) the effectiveness of adaptation to downstream tasks.
4.1 Training Efficiency
Experimental Setup
We utilize the PubMed Canese and Weis (2013) dataset, which comprises over 36 million records of biomedical literature. We extract around 400,000 tokens from the abstract of the literature to construct the training examples. We fine-tune the Gemma-2B (Mesnard et al., 2024), Llama2-7B and Llama2-13B (Touvron et al., 2023b) models using the generative pre-training objective with various efficient fine-tuning methods. We compare the results of full-tuning, freeze-tuning, GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the perplexity on the training examples to evaluate the efficiency of different methods. We also incorporate the perplexities of the pre-trained models as baselines. More experimental details can be found in Appendix D.1.
Results
The training efficiency results are presented in Table 3, where memory refers to the peak memory consumed during training, throughput is calculated as the number of tokens trained per second, and PPL represents the perplexity of the model on the training examples. Since full-tuning Llama2-13B lead to a memory overflow, the results are not recorded. We observe that QLoRA consistently has the lowest memory footprint because the pre-trained weights are represented in lower precision. LoRA exhibits higher throughput leveraging the optimization in LoRA layers by Unsloth. GaLore achieves lower PPL on large models while LoRA advantages on smaller ones.
4.2 Fine-Tuning on Downstream Tasks
Experimental Setup
To evaluate the effectiveness of different efficient fine-tuning methods, we compare the performance of various models after fine-tuning on downstream tasks. We construct the training set and test set using 2,000 examples and 1,000 examples from three representative text generation tasks, including CNN/DM (Nallapati et al., 2016), XSum (Narayan et al., 2018) and AdGen (Shao et al., 2019), respectively. We select several instruction-tuned models and fine-tune them following the sequence-to-sequence task using different fine-tuning methods. We compare the results of full-tuning (FT), GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the ROUGE score (Lin, 2004) on the test set of each task. We also incorporate the scores of the original instruction-tuned models as baselines. More experimental details can be found in Appendix D.2.
Results
The evaluation results on downstream tasks are shown in Table 4. We report the averaged scores over ROUGE-1, ROUGE-2 and ROUGE-L for each LLM and each dataset. Some results of Gemma-7B are not included in the table because the GaLore method is not applicable to this model. An interesting finding from the results is that LoRA and QLoRA achieve the best performance in most cases, except for the Llama2-7B and ChatGLM3-6B models on the CNN/DM and AdGen datasets. This phenomenon highlights the effectiveness of these efficient fine-tuning methods in adapting LLM models to specific tasks. Additionally, we observe that the Mistral-7B model performs better on English datasets while the Qwen1.5-7B model achieves higher scores on Chinese dataset. These results suggest that the performance of the fine-tuned models is also associated with their inherent capabilities on specific languages.
5 Conclusion and Future Work
In this paper, we demonstrate LlamaFactory, a unified framework for the efficient fine-tuning of LLMs. Through a modular design, we minimize dependencies between the models, datasets and training methods and provide an integrated approach to fine-tune over 100 LLMs with a diverse range of efficient fine-tuning techniques. Additionally, we offer a flexible web UI LlamaBoard, enabling customized fine-tuning and evaluation of LLMs without coding efforts. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks.
In the future, we will consistently keep LlamaFactory synchronous with the state-of-the-art models and efficient fine-tuning techniques. We also welcome contributions from the open-source community. In future versions, we will explore more advanced parallel training strategies and multimodal efficient fine-tuning of LLMs.
6 Broader Impact and Responsible Use
LlamaFactory has attracted a large number of individuals interested in LLMs to explore the possibility of fine-tuning their own models. This contributes significantly to the growth of the open-source community. It is gaining increasing attention and is being featured in Awesome Transformers [^3] https://github.com/huggingface/transformers/blob/v4.39.0/awesome-transformers.md as a representative of efficient fine-tuning frameworks for LLMs. We anticipate that practitioners build their LLMs upon our framework that bring benefits to society. Adherence to the model license is mandatory when using LlamaFactory for fine-tuning LLMs, thus preventing from any potential misuse.
Appendix A Related Work
In this section, we enumerate existing frameworks for fine-tuning LLMs, especially highlight those for efficient fine-tuning.
LLaMA-Adapter (Zhang et al., 2022) efficiently fine-tunes the Llama model (Touvron et al., 2023a) for instruction-following abilities using zero-init attention. FastChat (Zheng et al., 2023) is a framework focused on training and evaluating LLMs for chat completion purposes. LitGPT (AI, 2023) provides the implementation of generative models and supports various training methods. Open-Instruct (Wang et al., 2023d) is a fine-tuning framework that supports efficient fine-tuning. Colossal AI (Li et al., 2023b) takes advanced parallelism strategies for distributed training. LMFlow (Diao et al., 2023) is an extensible and efficient fine-tuning framework that supports decoder-only models, supporting both full-tuning and adapter tuning. GPT4All (Anand et al., 2023) allows LLMs to run on consumer devices, while also providing fine-tuning capabilities.
In contrast to the frameworks presented above, LlamaFactory supports a broader range of efficient fine-tuning methods. We compare our framework with existing work in Table 5.
Table 6: List of supported models.
Model Variant Organization
Llama (Touvron et al., 2023a) 7B/13B/33B/65B Meta AI
Llama 2 (Touvron et al., 2023b) 7B/13B/70B Meta AI
Baichuan Yang et al. (2023) 7B/13B Baichuan Inc
Baichuan2 (Yang et al., 2023) 7B/13B Baichuan Inc
BLOOM (Le Scao et al., 2022) 560M/3B/7.1B BigScience
BLOOMZ (Le Scao et al., 2022) 560M/3B/7.1B BigScience
ChatGLM2 (Du et al., 2022) 6B THUDM
ChatGLM3 (Du et al., 2022) 6B THUDM
ChineseLLaMA2 (Cui et al., 2023) 3B/7B/13B HFL
DeepSeek-LLM (Bi et al., 2024) 7B/67B DeepSeek
DeepSeek-Math (Shao et al., 2024) 7B DeepSeek
DeepSeek-MoE (Dai et al., 2024) 16B DeepSeek
DeepSeek-Coder (Guo et al., 2024) 6.7B/7B/33B DeepSeek
Falcon (Almazrouei et al., 2023) 7B/40B/180B TII
Gemma (Mesnard et al., 2024) 2B/7B Google
InternLM (Team, 2023) 7B/20B Shanghai AI Lab
InternLM2 (Team, 2023) 1.8B/7B/20B Shanghai AI Lab
Mistral (Jiang et al., 2023a) 7B Mistral AI
Mixtral (Jiang et al., 2023a) 8x7B Mistral AI
OLMo (Groeneveld et al., 2024) 1B/7B Allen AI
OpenChat (Wang et al., 2023b) 7B OpenChat
Orion (Chen et al., 2024a) 14B OrionStar
Phi-1.5 (Li et al., 2023c) 1.3B Microsoft
Phi-2 (Li et al., 2023c) 2.7B Microsoft
Qwen (Bai et al., 2023) 1.8B/7B/14B/72B Alibaba Cloud
Qwen1.5 (Bai et al., 2023) 1.8B/7B/14B/72B Alibaba Cloud
SOLAR (Kim et al., 2023) 10.7B Upstage AI
Skywork (Wei et al., 2023) 13B Skywork
StarCoder2 (Lozhkov et al., 2024) 3B/7B/15B BigCode
Vicuna1.5 (Zheng et al., 2023) 7B/13B LMSYS
Yi (Young et al., 2024) 6B/9B/34B 01.AI
Yuan2 (Wu et al., 2023) 2B/51B/102B IEIT
Zephyr (Tunstall et al., 2023) 7B Hugging Face H4
Appendix B Supported Models in LlamaFactory
We select the popular pre-trained models supported by LlamaFactory and list them in Table 6, ranging from dense models to sparse mixture-of-expert (MoE) models.
Table 7: Prompt of function calling in LlamaFactory.
You have access to the following tools:
> Tool Name: tool1
Tool Description: the usage of tool1
Tool Args:
- arg1 (dtype): the usage of arg1
Use the following format if using a tool:
Action: tool name (one of [tool1]).
Action Input: the input to the tool, in a JSON format representing the kwargs (e.g. ```{"input": "hello world", "num_beams": 5}```).
Appendix C Details of Designing Chat Template
Although Transformers 4.34.0 introduces the feature of chat templates, several models (such as ChatGLM and Qwen) have tokenizers that exhibit specific behaviours and cannot directly encode the special tokens. Therefore, we design a Formatter class to robustly convert the textual inputs to their embedding IDs. Specifically, we provide the EmptyFormatter, StringFormatter, FunctionFormatter and ToolFormatter. Furthermore, LlamaFactory supports fine-tuning models to obtain function calling capabilities. While the ReAct prompting Yao et al. (2023) is a popular choice for tool using, it is insufficient for nested tool parameters. We optimize the tool prompts and provided them in Table 7.
Appendix D Experiment Details
D.1 Training Efficiency
We adopt a learning rate of $10^{-5}$, a token batch size of $512$ to fine-tune these models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing to reduce the memory footprint. In freeze-tuning, we only fine-tune the last $3$ decoder layers of the model. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experiments are conducted on a single NVIDIA A100 40GB GPU. We enable flash attention in all experiments and Unsloth for LoRA and QLoRA experiments.
D.2 Fine-Tuning on Downstream Tasks
We briefly describe the datasets used for evaluating the effectiveness of different fine-tuning methods.
  • CNN/DM (Nallapati et al., 2016) This dataset is used for the summary generation task. English articles are gathered from CNN and Daily Mail. We utilize version 1.0.0, which comprises over 300k samples.
  • XSum (Narayan et al., 2018) This dataset is used for the abstractive summary generation task, which consists of English documents and one-sentence summaries, totaling over 200k samples.
  • AdGen (Shao et al., 2019) This dataset is a Chinese dataset that requires models to generate advertising text based on given keywords, totalling over 100k samples.
In this experiment, we adopt a learning rate of $10^{-5}$. We set batch size to $4$ while the maximum input length is $2048$. We fine-tune the models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing to reduce the memory footprint. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experimental results are carried out on the NVIDIA A100 40GB GPUs with flash attention enabled.